OpenStudy (anonymous):

-3[(w-6)^2+x]^2 W=5, X=6

OpenStudy (anonymous):

-3[(5-6)^2+6]^2 = -3[(-1)^2 + 6]^2 = -3 [1 + 6]^2 = -3 * 49 = -147

OpenStudy (anonymous):

dang